OpenStudy (anonymous):

Which is true? 1.The pressure in a fluid is constant. 2.Gravitational force decreases with increasing distance between two masses. 3.An object's weight varies depending on what planet the object is on. 4.All of the above.

OpenStudy (anonymous):

I haven't taken fluid mechanics yet, so I'm not sure about 1. However, that being said, I know enough to answer the question! For 2). we can refer to this equation\[\huge \text{F}_g=G\frac{m_1m_2}{r^2}\]And from this equation we can develop a relationship between the two parameters that we're asked in the question:\[\huge \text{F}_g \propto \frac{1}{r^2}\]What were to happen to the force if we increase the radius? For 3). we can refer to the equation for the weight of any object:\[\large \text{Weight}=\text{mass} \times \text{gravitational constant}=mg\]On Earth, the gravitational constant is \(9.81 m/s^2\), but on other planets, gravity is different! Assuming that the question only wants us to pick one correct answer, then we can immediately make the decision looking at these two.